OpenStudy (1645323):

Jessica and Franklin are volunteering to paint one of their community member’s fences for service hours. If Jessica could complete the job in 6 hours and Franklin could complete the job in 4.5 hours, how long would it take them to complete the job together, to the nearest minute? Select one: a. 2 hours and 34 minutes b. 2 hours and 44 minutes c. 2 hours and 57 minutes d. 3 hours and 16 minutes

OpenStudy (1645323):

@beccaboo333

OpenStudy (beccaboo333):

I'm not entirely sure how to do this.. Sorry :/ @ganeshie8

OpenStudy (1645323):

@whpalmer4 @apoorvk @thomaster

OpenStudy (1645323):

@iPwnBunnies @yellowlegoguy99

OpenStudy (apoorvk):

Means, Jessica completes (1/6) of the work alone in 1 hour, and Franklin completes (1/4.5) of the work in one hour. So, together, in ONE HOUR, together they can complete (1/6) + (1/4.5) of the work. Reciprocal of that would give the time they will need to complete the work together.

OpenStudy (1645323):

i got 1/27 @apoorvk

OpenStudy (apoorvk):

Incorrect. Re-calculate: \[\frac16 + \frac1{4.5}\]

OpenStudy (1645323):

oh 7/18

OpenStudy (1645323):

@ganeshie8 @ganeshie8 @ganeshie8 @ganeshie8

OpenStudy (1645323):

do you know what to do?

ganeshie8 (ganeshie8):

\(\large \dfrac{1}{6} + \dfrac{1}{4.5} = \dfrac{1}{x}\) \(\large \dfrac{7}{18} = \dfrac{1}{x}\) \(\large \dfrac{18}{7} = x\)

OpenStudy (1645323):

2.57 so that's 2 hours and 57 minutes?

OpenStudy (1645323):

nvm it is

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

its approximates 2 and HALF hours

ganeshie8 (ganeshie8):

how much is 2 and HALF hours ?

ganeshie8 (ganeshie8):

2.57 hours 2 hours + 0.57 hours 2 hours + 0.57*60 minutes 2 hours + 34 minutes

ganeshie8 (ganeshie8):

So, it is `2 hours and 34 minutes` okay ?

OpenStudy (1645323):

oh ok n ow I get it I didn't know you had to multiply .57 by 60

ganeshie8 (ganeshie8):

yes we need to cuz : \(0.57 \) hours is NOT same as \(57\) minutes

ganeshie8 (ganeshie8):

\(0.57\) is approximately HALF an hour... so you should get around \(30\) minutes